### v1

# Reinforcement Fine-Tuning of Large Language Models

[Uncaptioned image]: A General-Purpose and Unified Framework for
Reinforcement Fine-Tuning of Large Language Models
Xuchen Pan, Yanxi Chen, Yushuo Chen, Yuchang Sun, Daoyuan Chen,
Wenhao Zhang, Yuexiang Xie, Yilun Huang, Yilei Zhang, Dawei Gao,
Yaliang Li, Bolin Ding, Jingren Zhou
Alibaba Group
11footnotetext: Equal contribution.22footnotetext: Corresponding author. {chenyanxi.cyx, panxuchen.pxc, yaliang.li, bolin.ding}@alibaba-inc.com
Trinity-RFT is a general-purpose, flexible and scalable framework designed for reinforcement fine-tuning (RFT) of large language models. It is built with a decoupled design, consisting of (1) an RFT-core that unifies and generalizes synchronous/asynchronous, on-policy/off-policy, and online/offline modes of RFT, (2) seamless integration for agent-environment interaction with high efficiency and robustness, and (3) systematic data pipelines optimized for RFT. Trinity-RFT can be easily adapted for diverse application scenarios, and serves as a unified platform for exploring advanced reinforcement learning paradigms. This technical report outlines the vision, features, design and implementations of Trinity-RFT, accompanied by extensive examples demonstrating the utility and user-friendliness of the proposed framework.
Note: Trinity-RFT is currently under active development. This technical report corresponds to commit id f17db3d (May 23, 2025) of the GitHub repository, and will be continuously updated as the codebase evolves. Comments, suggestions and contributions are welcome!
Figure 1: The design of Trinity-RFT.
1 Vision of Trinity-RFT
Reinforcement learning (RL) has achieved remarkable success in the development of large language models (LLMs). Examples include aligning LLMs with human preferences via reinforcement learning from human feedback (RLHF) [18], and training long-CoT reasoning models via RL with rule-based rewards [3, 28]. However, such approaches are limited in their abilities to handle dynamic, real-world, and continuous learning.
Trinity-RFT envisions a future where AI agents learn by interacting directly with environments, collecting delayed or complex reward signals, and continuously refining their behavior through RL based on the collected experiences [26]. For example, imagine an AI scientist who designs an experiment, executes it, waits for feedback (while working on other tasks concurrently), and iteratively updates itself based on true environmental rewards when the experiment is finally finished. Trinity-RFT offers a path into this future by providing various useful features.
2 Key Features
Trinity-RFT is a general-purpose, flexible, scalable and user-friendly RL framework that can be adapted for diverse experimental or real-world scenarios. The following features of Trinity-RFT make it a promising solution for realizing the aforementioned vision:
  • Unified RFT modes & algorithm support. Trinity-RFT unifies and generalizes existing RFT methodologies into a flexible and configurable framework, supporting synchronous/asynchronous, on-policy/off-policy, and online/offline training, as well as hybrid modes that seamlessly combine them into a single learning process (e.g., incorporating expert trajectories to accelerate an online RL process [15, 36]). This is made possible partly by our decoupled design that allows rollout and training to be executed separately and scaled up independently on different machines, which will soon be introduced in the next section.
  • Agent-environment interaction as a first-class citizen. Trinity-RFT allows delayed rewards in multi-step/time-lagged feedback loops, and handles long-tailed latencies and environment/agent failures gracefully, which ensures efficiency and robustness in real-world scenarios with complex agent-environment interaction.
  • Systematic data pipelines optimized for RFT. These include converting raw datasets to task sets for RFT, cleaning/filtering/prioritizing experiences stored in the buffer, incorporating priors and additional reward signals from various sources into the collected experiences, synthesizing data for augmenting tasks and experiences, offering user interfaces for human-in-the-loop, among others.
3 Design and Implementations
The overall design of Trinity-RFT exhibits a trinity consisting of (1) RFT-core, (2) agent-environment interaction, and (3) data pipelines, which are illustrated in Figure 1 and elaborated in this section.
3.1 RFT-Core
RFT-core is the component of Trinity-RFT, highlighted at the center of Figure 1, where the core RFT process happens. Its design also exhibits a trinity, consisting of the explorer, trainer, and buffer.
  • The explorer, powered by a rollout model, takes a task as input and solves it by executing a workflow that specifies the logic of agent-environment interaction, thereby collecting experiences (including rollout trajectories, rewards, and other useful information) to be stored in the buffer.
  • The buffer stores experiences that can be generated by the explorer or by other sources, such as human experts. It also assists with fetching training samples for the trainer, and can be integrated with advanced sampling strategies and post-processing operations.
  • The trainer, backed by a policy model, samples batches of experiences from the buffer and updates the policy model via RL algorithms.
Our implementations allow the explorer and trainer to be deployed on separate machines. This decoupled design of RFT-core enables Trinity-RFT to support diverse RFT modes with great flexibility, for example:
  • In a synchronous mode, the explorer and trainer get launched simultaneously, work in close coordination, and synchronize their model weights once every sync_interval training steps. Within each synchronization period, the explorer continuously generates sync_interval batches of rollout experiences and stores them in the buffer, which are then retrieved and utilized by the trainer for updating the policy model. If sync_interval $=1$, this is an on-policy RL process, whereas if sync_interval $>1$, it becomes off-policy (akin to the mode adopted in [28]) and can be accelerated by pipeline parallelism between the explorer and trainer. This mode can be activated by setting the configuration parameter mode to both.
  • In a fully asynchronous mode, the explorer and trainer are mostly decoupled and can scale up independently. The explorer continuously generates rollout experiences and stores them in the buffer, while the trainer continuously samples experiences from the buffer and uses them for training the policy model. External experiences, e.g., those generated by expert models or humans, can be continuously incorporated into the buffer as well. The explorer synchronizes its model weights with the trainer once in a while, keeping the distribution of rollout experiences up to date. This mode can be activated by setting mode to explore/train and launching the explorer and trainer separately on different machines.
  • Trinity-RFT supports a benchmark mode that allows the user to evaluate one or multiple checkpoints on arbitrary benchmarks, after the RFT training process has finished. To activate this mode, the user simply needs to set mode to bench and specify the paths for the evaluation datasets in the configurations. This mode can be particularly useful for experimental purposes; for example, the user might want to try out different RFT techniques or configurations quickly (with limited evaluation on hold-out data) during training, identify which RFT trials have achieved stable convergence and high rewards, and then conduct more thorough evaluations only for the checkpoints of these successful trials.
  • In certain scenarios, the user would like to train the policy model without further exploration, using experiences that have already been collected and stored in the buffer. This train-only mode can be activated by setting the mode parameter to train and launching the trainer alone. Offline methods like Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) [20] can be regarded as special cases of such scenarios, both of which are natively supported by Trinity-RFT. For another example, consider an online RFT process that expands over a long period of time, where the explorer alone is launched during the daytime for serving human users and collecting experiences, while the trainer alone is launched at night for updating the policy model (which will be thoroughly validated and evaluated before it can be actually deployed as the rollout model for the next day).
Another benefit brought by the decoupled design is that explorers and trainers can operate across separate devices (e.g., multiple explorers on edge devices, one trainer on a high-end GPU cluster) and scale up independently. Indeed, we anticipate that for successful RL in real-world scenarios, distributed collection of massive rollout experiences could be one promising solution to resolving data scarcity.
Remark 1.
We take a system-algorithm co-design perspective in the development of Trinity-RFT, and aim to unify and generalize diverse RFT methodologies in this framework, including synchronous/asynchronous, on-policy/off-policy, and online/offline modes. RFT-core provides the necessary infrastructure for achieving this goal. While the main text of this technical report focuses on the system perspective, interested readers may refer to Appendix A for further discussions on the algorithm perspective of off-policy/asynchronous RL.
3.2 Agent-Environment Interaction
Numerous challenges arise when one tries to build an RFT framework that can efficiently and robustly handle real-world interaction between the LLM-powered agent and the environment. These include long-tailed latencies, agent/environment failures, and lagged reward signals, among others.
Trinity-RFT regards agent-environment interaction as a first-class citizen and incorporates various solutions to tackle these challenges, for example:
  • Built upon vLLM [10], our implementations of the LLM inference engine in Trinity-RFT support asynchronous and concurrent generation of rollout trajectories for multiple tasks. This helps mitigate the straggler effect caused by the long-tailed latencies in rollout generation and agent-environment interaction, thereby accelerating the RFT process. Load balancing among multiple LLM inference engines within one RFT training course is also taken care of, and would be one direction for further optimizing the utilization of computational resources.
  • Trinity-RFT incorporates various timeout/retry/skip mechanisms for fault tolerance and robustness, which ensure that continuous rollout generation would not be interrupted or blocked by individual failures in certain rounds of agent-environment interaction. This is crucial for stable and efficient learning in real-world scenarios, e.g., when the agent interacts with a large number of MCP services [12] that differ vastly in quality and availability.
  • Trinity-RFT is built to provide native support for asynchronous RFT modes, which allow great flexibility in the paces of the explorer and trainer. This can boost the overall efficiency of the RFT process, compared to synchronous modes where the slower one among the explorer and trainer can block the progress of the other and cause waste of computational resources.
  • For lagged reward signals, the trinity design of RFT-core offers a natural solution. As soon as the rollout trajectory (without reward values) has been generated, it is saved into the experience buffer, but marked as “not ready for training”. The explorer is now free from this task and may continue to collect experiences for other tasks. When the reward signals from the environment finally arrive, they are written to the buffer, and the corresponding experience is now marked as “ready for training”.
To adapt Trinity-RFT to a new downstream scenario, the user mainly needs to define and register a customized workflow (by inheriting the base class Workflow or MultiTurnWorkflow) where the logic of agent-environment interaction is implemented. Detailed examples will be demonstrated in Sections 4.1 and 4.2.
3.3 Data Pipelines
The data pipelines in Trinity-RFT aim to address fundamental challenges in RFT scenarios, such as managing heterogeneous data dynamics across interaction workflows, enabling delayed reward integration, and facilitating continuous data curation. Our solutions center on three core aspects: end-to-end data transformation, active data-reward co-shaping, and human-in-the-loop curation, each corresponding to key requirements identified in our development of RFT-core (Section 3.1).
3.3.1 End-to-end Data Transformation
To support the diverse RFT modes (e.g., synchronous or asynchronous) in Trinity-RFT, we establish a service-oriented data pipeline architecture as illustrated in Figure 2. It decouples data pipeline logic from procedure control to enable flexible RL-oriented data transformations with two key modules:
  • The Formatter Module unifies disparate data sources into RFT-compatible formats, providing convenient conversion between raw inputs (e.g., meta-prompts, domain-specific corpora, and QA pairs with tagged rewards) and structured RFT representations (e.g., via RftDatasetModel). For efficient RFT workloads, we utilize PostgreSQL-based persistent storage supporting both ExperienceModel for prioritized trajectories and DPODataModel for preference pairs, which is particularly beneficial for asynchronous RFT mode and distributed environments. Full schema details appear in Appendix B.1.
  • The Controller Module manages the complete data pipeline lifecycle through distributed server initialization, declarative configuration, and automated dataset persistence. It implements dynamic control mechanisms for asynchronous scenarios and protection against resource exhaustion, with configurable termination conditions based on compute quota or data quantity. This modular design enables Trinity-RFT to handle data transformations flexibly while maintaining consistency across different RFT modes. More service utilities are detailed in Appendix B.2.
The Formatter-Controller duality mirrors the explorer-trainer decoupling in RFT-core, enabling parallel data ingestion and model updating. This design also allows Trinity-RFT to handle delayed rewards through version-controlled experience updates while maintaining low-latency sampling for the trainer.
Figure 2: An illustration of data flow in Trinity-RFT.
3.3.2 Active Data-Reward Co-Shaping
To enable efficient and effective learning from complex and delayed reward signals, Trinity-RFT introduces fruitful mechanisms that actively shape both data and reward throughout the RFT process.
  • Trinity-RFT provides the Processor Module to support convenient and extensive data processing capabilities tailored to RFT, through dedicated bridge classes that link Trinity-RFT to over 100 operators from the Data-Juicer system [2]. This module offers efficient integration with the experience buffer, and provides composable building blocks for experience cleaning (e.g., length filters, duplication removal), safety alignment (e.g., toxicity detection, ethics checks), and preference data synthesis (e.g., critique-conditioned augmentation). Users can specify processing chains using either natural language or explicit configurations. These are accessible to both RFT users familiar with Data-Juicer and those who are not, thanks to transparency of the underlying system details. By treating Data-Juicer as a modular data processing operator pool rather than a central dependency, Trinity-RFT provides RL-specific abstractions and coherence, while benefiting from well-established data tools.
  • RL-oriented data shaping. RL is particularly useful in scenarios where optimization goals cannot be clearly defined, leading users to express vague or high-level requirements. In response, Trinity-RFT automates the translation of these objectives into optimized data processing pipelines. It reduces problem complexity through a structured hierarchy: Quality $\rightarrow$ Difficulty $\rightarrow$ Diversity $\rightarrow$ Quantity, with auto-selection of appropriate operators from Data-Juicer. Moreover, Trinity-RFT provides many composable and extensible interfaces for data shaping, aiming to seamlessly inject user-defined inductive biases. For instance, DataCleaner (a sub-module of Processor) supports multi-way cleaning with statistical filtering (e.g., entropy-based diversity [11]) and rule-based validation (e.g., boxed answer verification in math dataset [7]), while DataSynthesizer enables both autonomous generation and hybrid augmentation with controlled feedback metrics [35].
  • Active data optimization. As agents interact with environments, experiences collected and stored in the buffer can become large-scale and highly diverse. Using the right experiences for RL training is crucial for a cost-effective and stable learning process. To this end, Trinity-RFT allows users to flexibly apply multi-dimensional utility scoring to prioritize samples (e.g., 0.4 * difficulty + 0.3 * diversity + 0.1 * quality - 0.2 * frequency). Moreover, we introduce the DataActiveIterator module to enable (1) version-controlled experience reuse through ExperienceModel; (2) cross-task data lineage tracking via rich relationships derived from sample_id and rollout-model_id; and (3) incorporating human annotations into the rewards. These features will be demonstrated through concrete examples in Section 4 and Appendix B.3.
3.3.3 Human-AI Collaboration
In scenarios where human feedback is irreplaceable, Trinity-RFT establishes a bi-directional human-AI collaboration loop that provides first-class support for human annotations, based on Label Studio [30] and Data-Juicer’s HumanOPs.
  • Multi-stage annotation. Trinity-RFT implements configurable procedures combining automatic pre-screening and human verification. Typical stages include preference annotation (comparative assessment of model responses), quality auditing (human verification of automated cleaning/synthesis results), and cold-start bootstrapping (initial dataset curation through expert demonstrations).
  • Native asynchronism support. As the collection of human feedback is generally slower than AI/model feedback, we provide dedicated capabilities to handle both synchronous and asynchronous feedback modes, with configurable timeout and polling parameters. The feedback collaboration is based on an event-driven design, with automatic task creation upon data state changes, configurable notifications via email/Slack/webhook, and an atomic transaction model for annotation batches.
  • Customization. Different applications may involve humans in heterogeneous ways. We thus prioritize flexibility in both the interaction-interface and service levels. Examples include rich built-in interfaces that can be extended in a visualized style with XML-like tags provided by Label Studio, fine-grained quality scoring for reward shaping, free-form feedback attachment for dataset shaping, among others. Moreover, for easy deployment, we provide local Label Studio instance management with automatic environment setup via Docker/pip; optimized SDK interactions with batch request coalescing; unified logging across annotation tools and ML services; and concurrent annotation campaigns through priority-based task routing, while maintaining full data lineage preserved via LineageTracker.
The decoupled design of Trinity-RFT, or the presence of an experience buffer in particular, enables human feedback to participate in RL loops without breaking the asynchronous execution model. For instance, human-verified samples can be prioritized for training while fresh experiences are being collected, which is a critical capability for real-world deployment scenarios with mixed feedback sources. Further details for human-AI collaboration in Trinity-RFT will be illustrated in Section 4.4.
3.4 Performance Optimizations and User-Friendliness
Trinity-RFT has made numerous efforts in pursuit of efficiency, for example:
  • For multi-turn conversations and ReAct-style workflows [37], Trinity-RFT supports concatenating multiple rounds of agent-environment interaction compactly into a single sequence, with proper masking that indicates which tokens need to be incorporated into the training objective of RL algorithms.
  • For the synchronous RFT mode, Trinity-RFT accelerates the RFT process via pipeline parallelism between the explorer and trainer.
  • Trinity-RFT utilizes Ray [14] for distributed runtime, and NCCL communication primitives [16] (whenever feasible) for model weight synchronization between the explorer and trainer.
In addition, Trinity-RFT has been designed with user-friendliness as a top priority, for example:
  • We include a monitor (built upon Wandb [32] and TensorBoard [29]) that allows the user to conveniently track the progress of an RFT process, both quantitatively (e.g., via learning curves for rewards and other metrics) and qualitatively (e.g., via concrete examples of rollout trajectories generated at different RL steps). See Figure 3 for an example snapshot of the monitor.
  • We also implement a configuration manager, as shown in Figure 4, that allows the user to create configuration files conveniently via a front-end interface.
  • We provide Trinity-Studio, an all-in-one unified UI that allows the user to configure and run data inspection, data processing, RFT learning process, etc., all by clicking the mouse and filling forms, without writing any code. More functions will be integrated in the future. An example for developing with Trinity-Studio will be introduced in Section 4.5.
Figure 3: A snapshot of the monitor implemented in Trinity-RFT.
(a) The “beginner” mode.
(b) The “expert” mode.
Figure 4: Snapshots of the configuration manager.
4 Examples and Applications
Through some concrete examples, we demonstrate how to use Trinity-RFT for different purposes, and also exemplifies some concepts introduced in previous sections. Additional step-by-step tutorials can be found on the documentation website [^1] https://modelscope.github.io/Trinity-RFT, or the /examples folder of the GitHub repository [^2] https://github.com/modelscope/Trinity-RFT.
4.1 Single-Turn Workflow
In a simple yet common scenario, a user of Trinity-RFT would like to train an LLM for completing single-turn tasks, where the LLM generates one response to each input query. For this purpose, the user mainly needs to (1) define and register a single-turn workflow class (by inheriting the base class Workflow) tailored to the targeted tasks, and (2) specify the dataset of tasks (for training and/or evaluation) and the initial LLM, both compatible with HuggingFace [9] and ModelScope [13] formats.
Listing 1 gives a minimal example for implementing a single-turn workflow. Suppose that each task is specified by a <question, answer> tuple. The run() method of ExampleWorkflow calls the LLM once to generate a response for the question, calculates its reward, and returns an Experience instance that consists of the response itself, the reward value, and the log-probabilities of next-token prediction by the rollout model (which is necessary for certain RL algorithms, such as PPO [22] and GRPO [23]). Some built-in workflows have been implemented in Trinity-RFT, e.g., the MathWorkflow class for math-related tasks.
Trinity-RFT also allows more flexible and advanced usage by changing just a few configuration parameters, e.g., running supervised fine-tuning (SFT) as a warmup stage before the RFT process, or choosing between the synchronous and asynchronous modes of RFT (as described in Section 3.1).
1@WORKFLOWS.register_module("example_workflow")
2class ExampleWorkflow(Workflow):
3
4 def __init__(self, model: ModelWrapper, task: Task, **kwargs):
5 super().__init__(model, task, **kwargs)
6 self.question = task.raw_task.get("question")
7 self.answer = task.raw_task.get("answer")
8
9 def calculate_reward(self, response: str, truth: str) -> float:
10 return 1.0 if response == truth else 0.0
11
12 def run(self) -> List[Experience]:
13 response = self.model.chat(
14 [
15 {
16 "role": "user",
[…]
18 }
19 ],
20 n=self.task.rollout_args.repeat_times,
21 temperature=self.task.rollout_args.temperature,
22 )
23 reward: float = self.calculate_reward(response.response_text, self.answer)
24 return [
25 Experience(
26 tokens=response.tokens,
27 prompt_length=response.prompt_length,
28 reward=reward,
29 logprobs=response.logprobs,
30 )
31 ]
Listing 1: A minimal example for implementing a customized workflow.
4.2 Multi-Turn Workflow
In more advanced cases, the user would like to train an LLM-powered agent that solves multi-turn tasks by repeatedly interacting with the environment. With Trinity-RFT, achieving this is mostly as simple as in the single-turn case, except that the user needs to define and register a multi-turn workflow class (by inheriting the base class MultiTurnWorkflow), where the logic of agent-environment interaction is implemented. Listing 2 provides one such example for ALFWorld [25]. For training efficiency, the process_messages_to_experience() method concatenates multiple rounds of agent-environment interactions compactly into an Experience instance consisting of a single token sequence with proper masking, which can readily be consumed by standard RL algorithms like PPO and GRPO.
1@WORKFLOWS.register_module("alfworld_workflow")
2class AlfworldWorkflow(MultiTurnWorkflow):
3 """A workflow for the ALFWorld task."""
4
5 # ...
6
7 def generate_env_inference_samples(self, env, rollout_num) -> List[Experience]:
8 print("Generating env inference samples...")
9 experience_list = []
10 for i in range(rollout_num):
11 observation, info = env.reset()
12 final_reward = -0.1
13 memory = []
14 memory.append({"role": "system", "content": AlfWORLD_SYSTEM_PROMPT})
15 for r in range(self.max_env_steps):
16 format_obs = format_observation(observation)
17 memory = memory + [{"role": "user", "content": format_obs}]
18 response_text = self.get_model_response_text(memory)
19 memory.append({"role": "assistant", "content": response_text})
20 action = parse_action(response_text)
21 observation, reward, done, info = env.step(action)
22 if done:
23 final_reward = reward
24 break
25 experience = self.process_messages_to_experience(
26 memory, final_reward, {"env_rounds": r, "env_done": 1 if done else 0}
27 )
28 experience_list.append(experience)
29 # Close the env to save CPU memory
30 env.close()
31 return experience_list
32
33
34 def run(self) -> List[Experience]:
35 # ...
36 game_file_path = self.task_desc
37 rollout_n = self.repeat_times
38 # ...
39 env = create_environment(game_file_path)
40 return self.generate_env_inference_samples(env, rollout_n)
Listing 2: An implementation of a multi-turn workflow for ALFWorld [25].
4.3 Dataset Curation
We present an example to showcase the end-to-end data transformation (Section 3.3.1) and data-reward co-shaping (Section 3.3.2) capabilities of Trinity-RFT in a mathematical reasoning task. As shown in Listings 3 and 4, the user configures dataset paths and field mappings through a centralized DataConfig, then specifies natural-language processing instructions via dj_process_desc. The system will automatically translate these requirements into executable pipelines through three phases: (1) Processor applies LLM-powered difficulty scoring using Qwen-Max; (2) Controller optimizes quality-diversity tradeoffs through adaptive filtering thresholds and thereby prioritizes high-value training samples; (3) Formatter converts raw arithmetic problems into an RL-ready task set with specified reward functions and workflows.
This procedure exemplifies our active data optimization capability (Section 3.3.2) by combining automated quality assessments with configurable curation policies. Users can extend this pattern to many other domains through customization and adaptation: for example, users can define their own difficulty metrics and calculation rules in the parameters of Data-Juicer’s OPs. Besides, users can flexibly utilize another Synthesizer module and extend new Controller policies to support seamless transitions between different domain-specific reward shaping and data optimization priors.
1# Core dataset configuration
2class DataConfig:
3 """Centralized dataset configuration management"""
4
5 def __init__(self):
6 self.data = {
7 "dataset_path": "/path/to/gsm8k",
8 "format_config": {
9 "prompt_key": "question",
10 "response_key": "answer"
11 },
12 "dj_process_desc": "Compute difficulty scores for math questions",
13 "agent_model_name": "qwen-max"
14 }
15
16 def get_metadata(self) -> dict:
17 """Retrieve configuration metadata"""
18 return self.data
Listing 3: Core dataset configuration, which includes processing instructions as metadata.
1# Typical procedure that transforms raw dataset through multi-stage pipeline
2my_dataset = RftDataset(data_config) # Raw dataset auto-loading
3
4my_dataset.format([
5 BoxedMathAnswerFormatter(config), # Domain-specific formatting
6 RLHFFormatter(config) # RLHF format conversion
7 ])
8
9my_dataset.to_taskset(
10 reward_fn=AccuracyReward, # Reward shaping specification
11 workflow=MathWorkflow # Execution environment binding
12 )
Listing 4: Dataset transformation with domain-specific formatting and RLHF conversion.
4.4 Human-AI Collaborative Annotation
This example demonstrates the human-in-the-loop capability in Trinity-RFT for preference modeling. As illustrated in Listing 5 and Figure 5, our framework integrates Label Studio’s annotation interface with asynchronous data pipelines through four coordinated stages: (1) task generation: auto-creating annotation batches from model rollouts; (2) interactive labeling: providing UI for side-by-side response comparison; (3) quality control: enforcing inter-annotator agreement thresholds; and (4) versioned storage: tracking preference lineage in pre-defined fields like those in DPODataModel.
This pipeline reflects Trinity-RFT’s bi-directional collaboration feature (Section 3.3.3), backed by timeout-aware task polling and support of atomic batch commit. It enables hybrid procedures where initial AI pre-screening can reduce human workload in production deployments. Annotation activities can scale across distributed teams through event-driven task routing. The system’s flexibility benefits rapid adaptation to diverse annotation protocols, allowing developers to implement custom labeling interfaces through XML-based templates or integrate third-party annotation services via unified SDK endpoints. This capability underpins advanced use cases such as safety red-teaming datasets and online instruction tuning scenarios where human judgment remains irreplaceable for quality-critical decisions, particularly in human-centric sociocultural contexts where data quality, difficulty, and reward signals are difficult to verify logically.
1# Human annotation configuration
2class HumanAnnotationConfig:
3 """Preference annotation pipeline configuration"""
4
5 def __init__(self):
6 self.process = [
7 {
8 "human_preference_annotation_mapper": {
9 "wait_for_annotations": True, # Block until annotations complete
10 "timeout": 3600, # Maximum wait time in seconds
11 "prompt_key": "prompt", # Source field for prompts
12 "answer1_key": "answer1", # First candidate response
13 "chosen_key": "chosen" # Selected response key
14 }
15 }
16 ]
17
18 def get_pipeline(self) -> List[Dict]:
19 """Get annotation processing pipeline"""
20 return self.process
Listing 5: Configuration for human preference annotation.
Figure 5: An interactive interface for human preference annotation.
4.5 Trinity-Studio
Trinity-Studio provides visual interaction for the core capabilities of Trinity-RFT, designed to bridge the gap between system complexity and user accessibility. As shown in Figure LABEL:fig:trinity-studio:dashboard, its three integrated modules, “Training Portal”, pgAdmin”, and “Label Studio”, form a cohesive interface that directly supports the active data-reward co-shaping and human-AI collaboration capabilities established in Section 3.3.
  • “Training Portal” (Figure LABEL:fig:trinity-studio:training) implements configuration-to-execution procedures through declarative YAML editing with live validation and auto completion. This design aligns with Trinity-RFT’s decoupled architecture, enabling users to optionally specify dataset paths, processing and training parameters, meanwhile ensuring schema-aware validation against pre-defined data schema constraints (Section B.1). The live validation mechanism prevents misconfigurations that could disrupt the explorer-trainer coordination described in Section 3.1. Furthermore, the integration of runtime metrics with tools like Wandb/TensorBoard directly helps the active data optimization feature by surfacing signals such as difficulty distribution drifts and diversity metrics mentioned in Section 4.3. This transparency ensures that users can monitor how data curation strategies impact RFT performance in real time.
  • “pgAdmin” (Figure LABEL:fig:trinity-studio:pgadmin) reflects Trinity-RFT’s end-to-end data transformation capabilities by providing a visual panel for PostgreSQL-based storage (Section 3.3.1). This design benefits the versioned data lineage requirements of RFT, particularly for scenarios involving asynchronous training (Section 3.3.2). With intuitive SQL query builders, users can easily adjust schema, audit training experiences and human annotation batches with fine-grained precision. This capability is valuable for rapid validation of active learning policies by cross-referencing training outcomes with metadata (e.g., difficulty scores and staleness in asynchronous mode).
  • “Label Studio” page (Figure LABEL:fig:trinity-studio:label-studio) operationalizes Trinity-RFT’s bi-directional human-AI collaboration capability (Section 3.3.3). Utilizing the provided task polling and atomic batch commit mechanisms, users can annotate the data or experiences directly, allowing an asynchronous way to involve human feedback and to dynamically influence data curation.
By unifying these capabilities in a single UI, Trinity-Studio reduces the cognitive load of managing complex RFT procedures. For example, a researcher tuning a math reasoning task could use the Training Portal to adjust difficulty scoring parameters, view the resulting distribution shifts in the pgAdmin module, and then validate human annotators’ preferences in the Label Studio page. This end-to-end visibility is critical for debugging and iterating on RFT strategies, particularly when balancing tradeoffs between quality, diversity, and reward shaping. Moreover, this visual operation complements the programmatic APIs of Trinity-RFT while maintaining full compatibility with CLI procedures. We implement Trinity-Studio with Singe-Spa framework [27]. The modular architecture enables custom view development through JavaScript plugins and flexible extensions for general-purpose usage.
Figure 6: Snapshots of Trinity-Studio.
5 Conclusion
This report has provided an overview of Trinity-RFT, a general-purpose, flexible, scalable and user-friendly framework for reinforcement fine-tuning of large language models. Trinity-RFT offers a path into “the era of experience” [26], by supporting applications in diverse scenarios with complex agent-environment interaction, and serving as a unified platform for exploring advanced RL paradigms.
Acknowledgements
Trinity-RFT is built upon many excellent open-source projects, including but not limited to: verl [24] and PyTorch’s FSDP [5] for LLM training; vLLM [10] for LLM inference; Data-Juicer [2] for data processing functionalities; AgentScope [6] for agentic workflow; and Ray [14] for distributed systems. We have also drawn inspiration from RL/RFT projects like OpenRLHF [8], TRL [31], ChatLearn [1], and TinyZero [19].
Appendix A Off-Policy Reinforcement Learning Algorithms
For off-policy learning, one might simply adopt offline methods like supervised fine-tuning and direct preference optimization (DPO) [20], or their iterative variants where (some of) the training samples are generated online by the policy model itself and filtered by their reward values, e.g., rejection-sampling fine-tuning and iterative DPO. There also exist more advanced RL algorithms with dedicated design for off-policy or asynchronous scenarios [15, 21, 28, 4, 17, 33, 34, 36].
In the rest of this section, we present some of our findings about off-policy RL (with a focus on the bandit setting) through the lens of online policy mirror descent (OPMD). At the end of this section, we arrive at the surprising conclusion that the standard policy gradient, weighted by a coefficient and using the group mean reward as the baseline, can be a feasible direction for updating the policy even in off-policy settings (while the standard theory of policy gradient only holds for on-policy settings). This has been validated empirically in our exploratory experiments during the development of Trinity-RFT.
A.1 OPMD: Kimi’s Version
Below is a recap of the OPMD variant proposed in the technical report of Kimi k1.5 [28].
Analysis.
For a specific task/query $x$ and a reference policy $\pi_{\textsf{ref}}$, consider the following objective for training policy $\pi_{\bm{\theta}}$ at a particular iteration of the RL process:
$\max_{\bm{\theta}}\quad J(\bm{\theta};x,\pi_{\textsf{ref}})\coloneqq{\mathbb{E}}_{y\sim\pi_{\bm{\theta}}(\cdot|x)}[r(x,y)]-\tau\cdot D_{\textsf{KL}}\big{(}\pi_{\bm{\theta}}(\cdot|x)\|\pi_{\textsf{ref}}(\cdot|x)\big{)}.$
Note that $\pi_{\textsf{ref}}$ can be changing during the RL process. In [28], $\pi_{\textsf{ref}}$ is set to $\pi_{\bm{\theta}_{t}}$ at the $t$-th iteration, i.e., when updating the policy from $\bm{\theta}_{t}$ to $\bm{\theta}_{t+1}$.
The optimal policy $\pi^{\star}$ for this objective satisfies the following: for any response $y$,
$\pi^{\star}(y|x)=\frac{\pi_{\textsf{ref}}(y|x)e^{r(x,y)/\tau}}{Z}\propto\pi_{\textsf{ref}}(y|x)e^{r(x,y)/\tau},$ (1)
$\text{where}\quad Z\coloneqq Z(x,\pi_{\textsf{ref}})=\int\pi_{\textsf{ref}}(y^{\prime}|x)e^{r(x,y^{\prime})/\tau}\mathop{}\!\mathrm{d}y^{\prime}={\mathbb{E}}_{y^{\prime}\sim\pi_{\textsf{ref}}(\cdot|x)}[e^{r(x,y^{\prime})/\tau}].$ (2)
Taking logarithm of both sides of Eq. (1), we see that the optimal policy $\pi^{\star}$ must satisfy the following consistency condition:
$r(x,y)-\tau\cdot\log Z-\tau\cdot\big{(}\log\pi^{\star}(y|x)-\log\pi_{\textsf{ref}}(y|x)\big{)}=0.$
Algorithm.
Based on the above analysis, [28] proposes the following OPMD variant. For a query $x$, first sample $K$ rollouts $y_{1},\dots,y_{K}\sim\pi_{\textsf{ref}}(\cdot|x)$ from the reference policy, then define a surrogate loss as follows:
$\widehat{J}(\bm{\theta};x,\pi_{\textsf{ref}})\coloneqq\sum_{i\in[K]}\Big{(}r(x,y_{i})-\tau\cdot\log\widehat{Z}-\tau\cdot\big{(}\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\textsf{ref}}(y_{i}|x)\big{)}\Big{)}^{2},$
$\text{where}\quad\tau\cdot\log\widehat{Z}\coloneqq\tau\cdot\log\Big{(}\frac{1}{K}\sum_{i\in[K]}e^{{r(x,y_{i})}/{\tau}}\Big{)}.$
Although this is an off-policy method (since the rollout policy $\pi_{\textsf{ref}}$ is different from the policy $\pi_{\bm{\theta}}$ being updated), it is still limited because the rollouts have to be sampled from the particular policy $\pi_{\textsf{ref}}=\pi_{\bm{\theta}_{t}}$ for the $t$-th iteration of the RL process, as mentioned earlier. The reason for this limitation is the need of estimating $Z=Z(x,\pi_{\textsf{ref}})$ using samples from $\pi_{\textsf{ref}}(\cdot|x)$.
A.2 Pairwise OPMD
Analysis.
To eliminate the $Z$ term, we note that Eq. (1) is equivalent to the following:
$\forall\ y_{1}\text{ and }y_{2},\quad\frac{\pi^{\star}(y_{1}|x)}{\pi^{\star}(y_{2}|x)}=\frac{\pi_{\textsf{ref}}(y_{1}|x)}{\pi_{\textsf{ref}}(y_{2}|x)}e^{\big{(}r(x,y_{1})-r(x,y_{2})\big{)}/\tau}.$
Taking logarithm of both sides, we have
$\log\pi^{\star}(y_{1}|x)-\log\pi^{\star}(y_{2}|x)=\log\pi_{\textsf{ref}}(y_{1}|x)-\log\pi_{\textsf{ref}}(y_{2}|x)+\frac{r(x,y_{1})-r(x,y_{2})}{\tau},$
or equivalently,
$r(x,y_{1})-\tau\cdot\big{(}\log\pi^{\star}(y_{1}|x)-\log\pi_{\textsf{ref}}(y_{1}|x)\big{)}=r(x,y_{2})-\tau\cdot\big{(}\log\pi^{\star}(y_{2}|x)-\log\pi_{\textsf{ref}}(y_{2}|x)\big{)}.$
Note that this holds true for a pair of arbitrary responses $y_{1}$ and $y_{2}$.
Algorithm.
For a query $x$ and $K$ arbitrary responses $y_{1},\dots,y_{K}$, we define the following surrogate loss:
$\widehat{J}(\bm{\theta};x,\pi_{\textsf{ref}})\coloneqq\sum_{1\leq i<j\leq K}\big{(}a_{i}-a_{j}\big{)}^{2},$
$\text{where}\quad a_{i}\coloneqq r(x,y_{i})-\tau\cdot\big{(}\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\textsf{ref}}(y_{i}|x)\big{)},\quad i\in[K].$
Here $\pi_{\textsf{ref}}$ can be any reference policy for KL regularization, regardless of how $y_{1},\dots,y_{K}$ were sampled. While this is a fully off-policy RL method, it has its own limitation: to run this algorithm, we should make sure that multiple (at least 2) rollouts for the same task are included within one micro-batch (whose size is typically much smaller that of a batch or mini-batch), which adds to infrastructure complexity.
Remark 2.
In the special case of $K=2$, the above method, termed as “pairwise OPMD”, turns out to be the same as “contrastive policy gradient” proposed in [4], albeit with a simpler and more intuitive derivation.
A.3 OPMD: an Embarrassingly Simple Variant
Analysis.
Consider the $t$-th iteration of the RL process, i.e., updating from $\bm{\theta}_{t}$ to $\bm{\theta}_{t+1}$, and use $\pi_{\textsf{ref}}=\pi_{\bm{\theta}_{t}}$ as the reference policy. For a specific task/query $x$, recall from Section A.1 the original objective:
$\max_{\bm{\theta}}\quad J(\bm{\theta};x,\pi_{\bm{\theta}_{t}})\coloneqq{\mathbb{E}}_{y\sim\pi_{\bm{\theta}}(\cdot|x)}[r(x,y)]-\tau\cdot D_{\textsf{KL}}\big{(}\pi_{\bm{\theta}}(\cdot|x)\|\pi_{\bm{\theta}_{t}}(\cdot|x)\big{)}.$
We leverage the analysis in Section A.2, and take a closer look at the following pairwise loss for $a_{i}$ and $a_{j}$, normalized by $1/(1+\tau)^{2}$ to make the loss scale invariant to the hyperparameter $\tau$:
$\frac{(a_{i}-a_{j})^{2}}{(1+\tau)^{2}}$ $=\frac{1}{(1+\tau)^{2}}\bigg{[}\Big{(}r(x,y_{i})-r(x,y_{j})\Big{)}$
$\quad\quad-\tau\cdot\Big{(}\big{(}\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\bm{\theta}_{t}}(y_{i}|x)\big{)}-\big{(}\log\pi_{\bm{\theta}}(y_{j}|x)-\log\pi_{\bm{\theta}_{t}}(y_{j}|x)\big{)}\Big{)}\bigg{]}^{2}.$
The trick here is that, if we only intend to take one gradient step of this loss at $\bm{\theta}=\bm{\theta}_{t}$, then the value of $(\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\bm{\theta}_{t}}(y_{i}|x))-(\log\pi_{\bm{\theta}}(y_{j}|x)-\log\pi_{\bm{\theta}_{t}}(y_{j}|x))$ is simply zero. As a result,
$\nabla_{\bm{\theta}}\frac{(a_{i}-a_{j})^{2}}{(1+\tau)^{2}}\Big{|}_{\bm{\theta}_{t}}=\frac{-2\tau}{(1+\tau)^{2}}\Big{(}r(x,y_{i})-r(x,y_{j})\Big{)}\Big{(}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}}-\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{j}|x)|_{\bm{\theta}_{t}}\Big{)},$
and thus
$\nabla_{\bm{\theta}}\sum_{1\leq i<j\leq K}\frac{(a_{i}-a_{j})^{2}}{(1+\tau)^{2}}\Big{|}_{\bm{\theta}_{t}}$
$\quad=\sum_{1\leq i<j\leq K}\frac{-2\tau}{(1+\tau)^{2}}\Big{(}r(x,y_{i})-r(x,y_{j})\Big{)}\Big{(}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}}-\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{j}|x)|_{\bm{\theta}_{t}}\Big{)}$
$\quad=\sum_{1\leq i<j\leq K}\frac{-2\tau}{(1+\tau)^{2}}\bigg{(}\big{(}r(x,y_{i})-r(x,y_{j})\big{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}}+\big{(}r(x,y_{j})-r(x,y_{i})\big{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{j}|x)|_{\bm{\theta}_{t}}\bigg{)}$
$\quad=\frac{-2\tau}{(1+\tau)^{2}}\sum_{1\leq i\leq K}\sum_{1\leq j\leq K}\big{(}r(x,y_{i})-r(x,y_{j})\big{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}}$
$\quad=\frac{-2\tau}{(1+\tau)^{2}}\sum_{1\leq i\leq K}K\cdot\big{(}r(x,y_{i})-\overline{r}(x)\big{)}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x)|_{\bm{\theta}_{t}},$
where $\overline{r}(x)\coloneqq\frac{1}{K}\sum_{j\in[K]}r(x,y_{j})$ in the last line.
Algorithm.
To this end, we update from $\bm{\theta}_{t}$ to $\bm{\theta}_{t+1}$ by taking one gradient step of the following surrogate loss, where we simplify the constant factor from $2\tau/(1+\tau)^{2}$ to $1/(1+\tau)$ and also drop the $K$ factor:
$\min_{\bm{\theta}}\quad\widehat{J}(\bm{\theta};x)\coloneqq-\frac{1}{1+\tau}\sum_{1\leq i\leq K}\big{(}r(x,y_{i})-\overline{r}(x)\big{)}\log\pi_{\bm{\theta}}(y_{i}|x).$
This is exactly the standard policy gradient using the group mean reward as the baseline, but derived differently with an off-policy interpretation. The hyperparameter $\tau\geq 0$ controls the size of each policy update. A larger $\tau$ is effectively the same as a smaller learning rate.
As a heuristic, we simply add a regularization term (denoted by $g$) to the above objective when additional regularization with respect to a fixed policy, e.g., a SFT model $\pi_{\textsf{sft}}$, is desired:
$\min_{\bm{\theta}}\quad\widehat{J}(\bm{\theta};x)\coloneqq-\frac{1}{1+\tau}\sum_{1\leq i\leq K}\big{(}r(x,y_{i})-\overline{r}(x)\big{)}\log\pi_{\bm{\theta}}(y_{i}|x)+\beta\cdot g\big{(}\pi_{\bm{\theta}},\pi_{\textsf{sft}};x,y_{1},\dots,y_{K}\big{)}.$
Appendix B Additional Details for Data Pipelines
B.1 Illustration of Typical Dataset Classes and Schemas
The dataset shaping process described in Section 3.3 is supported by PostgreSQL persistence with comprehensive schema designs; one example can be found in Listing 6. Key model classes include RftDatasetModel (core dataset storage with lineage tracking), ExperienceModel (which stores serialized experiences with priority scoring), DPODataModel and SFTDataModel (specialized storage for DPO and SFT, respectively).
1-- RftDataset table structure
2CREATE TABLE rft_dataset (
3 id SERIAL PRIMARY KEY, -- Unique record identifier
4 consumed_cnt INTEGER DEFAULT 0, -- Usage counter for sampling
5 prompt TEXT, -- Input instruction
6 response TEXT, -- Model-generated response
7 reward FLOAT, -- Training signal value
8 quality_score FLOAT DEFAULT 0.0, -- Quality assessment metric
9 difficulty_score FLOAT DEFAULT 0.0, -- Complexity estimation
10 diversity_score FLOAT DEFAULT 0.0, -- Response uniqueness measure
11 priority FLOAT DEFAULT 0.0 -- Sampling preference weight
12);
Listing 6: Schema definition for RftDataset PostgreSQL table.
B.2 Illustration of Data Processing Service
The architecture presented in Section 3.3 integrates Data-Juicer capabilities through service-oriented design. Typically, the DJ-service lifecycle covers three stages: initialization via scripts/start_servers.py; configuration through dj_config_path or natural language (dj_process_desc); and execution handling dataset loading, cleaning, priority scoring, and database export. Moreover, we provide DJExecutor to bridge Data-Juicer capabilities through config translation and execution sandboxing, and dynamic termination via _check_clean_targets() with melt protection.
B.3 An Example of Active Data Optimization Algorithms
The DataActiveIterator serves as the core engine for data-model co-evolution with multi-dimensional utility scoring. It provides extensible interfaces for flexible training prior injection, e.g., as demonstrated in Algorithm 1.
Algorithm 1 Multi-Dimensional Sample Selection
0:  Dataset $D$, weights $w_{q},w_{d},w_{k},w_{f}$
0:  Selected subset $D^{\prime\prime}$
1:  $(q,d,k)\leftarrow\text{DataCleaner}.\text{Process}(D)$ Compute quality ($q$), diversity ($d$), difficulty ($k$)
2:  $U_{i}\leftarrow w_{q}q_{i}+w_{d}d_{i}+w_{k}k_{i}-w_{f}f_{i}$ Combine $q$, $d$, $k$, frequency ($f$)
3:  $D^{\prime}\leftarrow\{s_{i}\in D\mid U_{i}>\theta_{\text{min}}\}$ Threshold filtering
4:  $D^{\prime\prime}\leftarrow\text{Top-K}(D^{\prime},K)$ Priority selection
5:  $\text{ExperienceCache}.\text{Update}(D^{\prime\prime})$ Update version-controlled cache
6:  $\text{LineageTracker}.\text{Record}(D^{\prime\prime})$ Track data provenance
7:  return $D^{\prime\prime}$